OpenStudy (anonymous):

this is one of the last 2 questions i need will someone plz help me!!! Find the polynomial f(x) that has the roots of –3, 5 of multiplicity 2. Part 2: Explain how you would verify the zeros of f(x).

OpenStudy (anonymous):

and i need it like explained :/

OpenStudy (anonymous):

(x+3)(x-5)(x-2)^2=0

OpenStudy (anonymous):

is the answer?

OpenStudy (anonymous):

i need to know how to do it so i can do the others..

OpenStudy (anonymous):

well, if x=-3 the x+3=0 if x=5 then x-5=0 if x=2 with multiplicty 2 (x-2)^2=0